OpenStudy (anonymous):

solve equation by square root property y(sqrt2) - 250 = 0

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y^2 - 250 = 0 y^2 = 250 I'll let you finish up

OpenStudy (anonymous):

@jim_thompson5910 125?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you have to take the square root of both sides and simplify

jimthompson5910 (jim_thompson5910):

don't forget about the plus/minus

OpenStudy (anonymous):

do i do sq root of 2 and sq root of 250?

jimthompson5910 (jim_thompson5910):

y = sqrt(250) y = ???

OpenStudy (anonymous):

62500

jimthompson5910 (jim_thompson5910):

square root, not square

jimthompson5910 (jim_thompson5910):

250^2 = 62500 but the square root, or sqrt, is the opposite

OpenStudy (anonymous):

oh u divide?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

square root

OpenStudy (anonymous):

so 250 x 250?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

idk how you're getting that

jimthompson5910 (jim_thompson5910):

square root is not the same as squaring a number

OpenStudy (anonymous):

i never knew that

OpenStudy (anonymous):

so what do i do?

jimthompson5910 (jim_thompson5910):

example 6 squared = 6^2 = 6*6 = 36 going in reverse the square root of 36 = sqrt(36) = 6

jimthompson5910 (jim_thompson5910):

see the difference?

OpenStudy (anonymous):

so 250 x 2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

y = sqrt(250) how would you simplify that

OpenStudy (anonymous):

250 x 250?

jimthompson5910 (jim_thompson5910):

which number in the list below 1, 4, 9, 16, 25, 36, 49, 64, 81, 100 is a factor of 250 ?

OpenStudy (anonymous):

25 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so 250 = 25*10

jimthompson5910 (jim_thompson5910):

which means y = sqrt(250) y = sqrt(25*10) y = sqrt(25)*sqrt(10) y = 5*sqrt(10)

jimthompson5910 (jim_thompson5910):

so the two solutions are y = 5*sqrt(10) or y = -5*sqrt(10)

jimthompson5910 (jim_thompson5910):

don't forget about the plus/minus

OpenStudy (anonymous):

how do u know which one to choose?

jimthompson5910 (jim_thompson5910):

both are solutions

jimthompson5910 (jim_thompson5910):

since both satisfy y^2 - 250 = 0

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw